The Queen is placing the pen in an area with 500 square meters of space. She wants the pen to take up no more than 1/3
of the space available.

Which equation correctly represents this situation?
1
Expert's answer
2020-12-16T20:07:02-0500

Solution:

"area\,of\,space\,available, A = 500m^2 \\
area\,of\,space\,take\, by\,pen, A_P = \frac{500}{3}m^2 \\
therefore, \\
0 \leq A_p \leq A \\
\Rightarrow 0 \leq \frac{500}{3} \leq 500 \,\,\,\,\,\,\,\, \{ the\, required\,equation \}"